0.4(x-6)=0.7(2x+3)+0.5
Answer by ethai18(3)

0.4(x-6)=0.7(2x+3)+0.5
0.4x-2.4=1.4x+2.1+0.5
0.4x-2.4=1.4x+2.6
-x=5
x=-5